OpenStudy (anonymous):

9/10 greater than or less or equal 7/8

OpenStudy (jamesj):

Well, 9/10 = 1 - 1/10 and 7/8 = 1 - 1/8 Now, which is smaller 1/10 or 1/8 ?

OpenStudy (anonymous):

1/8

OpenStudy (jamesj):

Would you rather have 1/8 of a pie or 1/10 of a pie?

OpenStudy (anonymous):

1/0

OpenStudy (jamesj):

No. 1/10 is smaller than 1/8.

OpenStudy (anonymous):

Oh :/

OpenStudy (jamesj):

1/2 is more than 1/3. 1/4 is less than 1/3. 1/5 is more than 1/6.

OpenStudy (jamesj):

So tell me again, which is larger: 1/8 or 1/10 ?

OpenStudy (anonymous):

1/8

OpenStudy (jamesj):

Right. Hence which is larger: 1 - 1/10 or 1 - 1/8 ?

OpenStudy (anonymous):

1/10th of a pie if you're overweight ^^

OpenStudy (unklerhaukus):